Assume you use a revolutionary green car, whose fuel consumption per hour is proportional to the square root the car’ speed. You have to travel 750 km from Windhoek to a village at a constant speed, in a dust storm with a strong headwind of 25 km per hour. Assume that the minimum speed allowed on that road is 20km/h and the maximum speed (due to the storm) is 80 km/h. At what speed should you travel to minimize your fuel consumption. [20]
1
Expert's answer
2020-05-25T18:32:52-0400
"20\le v-25 \le 80"

"20+25 \le v \le80+25"

"45 \le v \le105"

Ansver: 45 km/h